# Benchmarking Time Series Foundation Models for Short-Term Household Electricity Load Forecasting

### v1

@@ 3.3. Exploratory Data Analysis (EDA)
Figure 5. Median hourly energy consumption per month
[…]
Median hourly energy consumption per month
Figure 5 shows the median consumption per hour in the different months of the year. There is a clear reduction in consumption from spring to fall in comparison to the winter months. This change is very apparent in the Southern Germany and East Midlands datasets, and less visible in Lower Saxony and Lothian.
It is important to consider that the datasets contain distinct numbers of households and different household characteristics, such as size, number of occupants, energy needs, the presence or absence of PV installations, and sometimes the use of heat pumps for heating. This last point could explain why the consumption in some datasets differs from the rest and is more nuanced by the seasons. The focus of this study is not to explain the causes of the differences in the datasets, but to use them to compare the models. Such heterogeneous datasets represent a realistic view of energy consumption modeling challenges, where the DSOs have insights into the energy demand causes in individual households.
The datasets have individual missing values between two measurements or, in the case of the Lower Saxony and East Midlands datasets, up to weeks-long gaps. The reasons given in the datasets’ descriptions are technical failures of the data logger, Internet outages, conversion work (Schlemminger et al., 2022), failed radio transmissions, and problems in daylight savings time transitions (Open Power System Data, 2020).
@@
  • the Lower Saxony and East Midlands dataset has long gaps in the measurements.
@@
Table 5. Comparison of models and their architectures and implementations
@@
3.4. Data Preprocessing
3.4.1. Handling Missing Values
3.4.2. Training and Test Split
Figure 6. Train-Test split. Time spans of each time series
3.5. Evaluated Models
A condition for the selection of foundation models was the disclosure of the training data and the provision of a pre-trained model as discussed in Chapter 3.2. At the time of writing, Chronos (Ansari et al., 2024), TimesFM (Das et al., 2024) and LagLlama (Rasul et al., 2024) fulfill both conditions. The foundation models are not be fine-tuned on the household datasets. Only inference is used on the unseen data for achieving zero-shot predictions (Xian et al., 2019). The Chronos model is implemented via Github as suggested by Ansari et al. (2024) and the model “chronos-t5-large” is used. The same applies for TimesFM (Das et al., 2024) and LagLlama (Rasul et al., 2024) by implementing them from the official Github Repo using the provided model weights. For LagLlama, we activate RoPe scaling as suggested in the zero-shot tutorial for input sizes longer than the context length of 32, because most of the tested input sizes are above that threshold. As the main TFS competitors for the foundation models, we used time series models with a similar Transformer-based architecture. Namely, we chose the original Transformer Encoder-Decoder implementation (”VanillaTransformer”) (Vaswani et al., 2017), and PatchTST (Nie et al., 2023), which represents the current SOTA for household energy STLF (see Table 1). Furthermore, we chose the relatively new iTransfomer (Liu et al., 2024) and Temporal Fusion Transformer (TFT) (Lim et al., 2021), which we identified in our literature review. The TFS Transformers are trained univariate, which can be a limiting factor for the TFS Transformers to varying degrees. However, this approach allows for the most direct comparison with the foundation models. For implementing the TFS Transformers, we used the NeuralForecast library, which is referring its model implementations to the corresponding papers.
As a baseline, we take the SeasonalAverage (Hyndman and Athanasopoulos, 2018) with a seasonality of 24 as we observed a high autocorrelation in the data with lags of 24. The SeasonalAverage is calculated per time series with the Pandas library with an hourly average over the input size as forecasts for the horizon.
This study’s primary objective is to assess the different approaches’ general ability to adapt to new data. We vary the input size for all models to identify the influence of data context for pattern recognition. We also refrain from extensive hyperparameter tuning. We gathered over 2 million data points, which, in combination with compute-intensive TFS Transformer training, would require a substantial amount of compute resources for hyperparameter tuning. Thus, we use the default parameters of the models as proposed in their implementation.
A complete overview of the used models and hyperlinks to their implementation can be found in Table 5.
3.6. Training and Evaluation
We follow the approach of a time series cross-validation (Bergmeir and Benítez, 2012) with a calibration window. The TFS Transformers are trained on a fixed window size (365 days) of the most recent observations up to the split date. After the initial forecast, both the split date and the calibration window are moved forward by the duration of the horizon. Consequently, the training data now encompasses the most recent 365 days, starting from this updated split date. The model is retrained from scratch on the updated calibration window, generating new forecasts for the next horizon period.
@@ 3.6. Training and Evaluation
Following the insights that the global is the best training approach (Hertel et al., 2023), we train the TFS Transformers on all the datasets’ training splits together.
@@
3.7. Challenges in Evaluating Foundation Models
While we prevent information leakage of global temporal patterns in the globally trained TFS Transformer models, there remains a special challenge in evaluating time series foundation models like Chronos (Ansari et al., 2024), TimesFM(Das et al., 2024) or LagLlama (Rasul et al., 2024). The pre-trained models might have learned global temporal patterns (e.g., Covid-19 (Prabowo et al., 2023), geopolitical crises) due to the massive amount of data used during training. Any evaluation on a dataset that is in the same date range as the training data of the foundation model could be potentially affected by such information leakage. Even cross-domain influences are possible, e.g., as weather affects household energy consumption (Kang and Reiner, 2022), a foundation model trained on historical weather data could theoretically lead to information leakage on a household energy dataset in the same time period.
@@ 3.7. Challenges in Evaluating Foundation Models
3.8. Metrics
The prediction results and the true results are scaled by subtracting the means and dividing by the standard deviation to be able to compare the time series with different demand ranges, for instance, given by the sizes of the houses and the number of occupants. The focus of the evaluation is the relative performance between models.
@@ 3.8. Metrics
(1) $MAE_{h}=\frac{1}{h}\sum_{h=1}^{h}\left(\frac{1}{n}\sum_{i=1}^{n}|y_{h,i}-\hat{y}_{h,i}|\right)$
@@
The same logic is also used to calculate the Mean Squared Error households (MSEh).
@@
Based on the various occurrences of zero or near-zero values, we discard other metrics, such as the mean percentage error (MAPE) and symmetric MAPE (SMAPE), as they tend to increase drastically with values near to zero (Hyndman and Koehler, 2006).
4. Results
Table 6 shows the MAEh and MSEh for all models across the different datasets. It can be seen that TimesFM achieved the overall best MAEh, closely followed by PatchTST. Only for the Lothian dataset, PatchTST outperformed TimesFM in terms of MSEh. The worst-performing model was LagLlama, which could not overperform the SeasonalAverage baseline.
@@ 4. Results
Table 6. Model´s MAEh and MSEh scores accross all datasets
Lothian Lower Saxony East Midlands Southern Germany
Model MAEh MSEh MAEh MSEh MAEh MSEh MAEh MSEh
TimesFM (F) 0.5139 0.9975 0.4909 0.9291 0.5307 0.8860 0.4781 0.7229
PatchTST (TFS) 0.5156 0.9685 0.4940 0.9550 0.5356 0.8992 0.4996 0.7601
Chronos (F) 0.5367 1.1841 0.5281 1.1673 0.5762 1.1226 0.5214 0.9243
VanillaTransformer (TFS) 0.5510 1.0812 0.5590 1.1041 0.6073 1.1240 0.5773 0.9037
TFT (TFS) 0.5767 1.1253 0.5797 1.1470 0.6349 1.1715 0.6152 0.9718
iTransformer (TFS) 0.5890 1.1331 0.5875 1.1032 0.6477 1.1169 0.5970 0.9188
SeasonalAverage (B) 0.6090 1.2082 0.5715 1.1024 0.5774 0.9814 0.5387 0.8532
LagLLama (F) 0.6463 1.3659 0.6705 1.3884 0.7434 1.4467 0.6443 1.1464
Table 7 shows the expanded MAEh by model, input size, and horizon. PatchTST exhibits the lowest forecasting error for the 24-hour input size for all horizons, followed by TimesFM and VanillaTransformer. TimesFM performed best for input sizes of 96 and 168 hours across all horizons, followed by PatchTST and Chronos. LagLlama, TFT and iTransformer performed poorly for 96 and 168 horizons, not being able to outperform the SeasonalAverage baseline.
@@
Table 7. MAEh results for every input size and horizon for all datasets
Model Input Horizon
@@
PatchTST (TFS) 24 0.5116 0.5276 0.5404
TimesFM (F) 24 0.5252 0.5379 0.5703
VanillaTransformer (TFS) 24 0.5286 0.5590 0.5723
Chronos (F) 24 0.5477 0.5533 0.5656
TFT (TFS) 24 0.5634 0.5748 0.5895
LagLLama (F) 24 0.5990 0.7781 0.8168
iTransformer (TFS) 24 0.6180 0.5958 0.5951
SeasonalAverage (B) 24 0.6351 0.6701 0.7119
TimesFM (F) 96 0.4921 0.5041 0.5113
PatchTST (TFS) 96 0.5013 0.5116 0.5153
Chronos (F) 96 0.5198 0.5347 0.5458
VanillaTransformer (TFS) 96 0.5252 0.5677 0.5640
SeasonalAverage (B) 96 0.5728 0.5814 0.5759
LagLLama (F) 96 0.5768 0.6605 0.6832
TFT (TFS) 96 0.5800 0.5900 0.5668
iTransformer (TFS) 96 0.6082 0.5878 0.5714
TimesFM (F) 168 0.4751 0.4899 0.4956
PatchTST (TFS) 168 0.4977 0.5064 0.5138
Chronos (F) 168 0.5053 0.5241 0.5367
VanillaTransformer (TFS) 168 0.5347 0.5650 0.5667
SeasonalAverage (B) 168 0.5478 0.5548 0.5598
LagLLama (F) 168 0.5527 0.6009 0.6202
TFT (TFS) 168 0.5836 0.5952 0.5867
iTransformer (TFS) 168 0.6060 0.5713 0.5762
Figure 7 illustrated the percentage difference from the SeasonalAverage baseline for every input size and horizon across all datasets. White denotes a percentage difference of 0, while worse and better performance in relation to the baseline is shown with shades of red and blue, respectively. The best performance compared to the baseline (i.e., 24.1 % improvement) is the one of PatchTST with an input size of 24 and a horizon of 168. The worst performance (i.e., 18.6% decrease) comes from LagLlama with an input size of 96 and a horizon of 168.
Figure 7. Percentage Difference between the models and the baseline for all datasets.
@@
Overall, the models’ mean error increases with longer horizons and decreases with longer input sizes. The SeasonalAverage’s performance also increases with the input size, making it a strong baseline. Furthermore, the absolute error diminishes with the input size while the percentage difference to the baseline decreases.
LagLlama tends to demonstrate better performance when provided with a longer input size, although it is unable to match the overall performance of the other forecasting models. For almost all models, the error also increases as the horizon rises.
5. Discussion
Our experiments on four datasets, namely Lower Saxony, Southern Germany, Lothian, and East Midlands, show that while the best foundation model, TimesFM, slightly outperforms the best TFS Transformer, PatchTST, not all foundation models perform equally well.

### v2

@@ 3.2 Exploratory Data Analysis (EDA)
Figure 5 shows the median consumption per hour in the different months of the year. There is a clear reduction in consumption from spring to fall in comparison to the winter months. This change is very apparent in the Southern Germany and REFIT datasets, and less visible in the Lower Saxony and IDEAL dataset.
It is important to consider that the datasets contain distinct numbers of households and different household characteristics, such as size, number of occupants, energy needs, the presence or absence of PV installations, and sometimes the use of heat pumps for heating. This last point could explain why the consumption in some datasets differs from the rest and is more nuanced by the seasons. The focus of this study is not to explain the causes of the differences in the datasets, but to use them to compare the models. Such heterogeneous datasets represent a realistic view of energy consumption modeling challenges, where the DSOs have no insights into the energy demand causes in individual households.
The datasets have individual missing values between two measurements or, in the case of the Lower Saxony and REFIT datasets, up to weeks-long gaps. The reasons given in the datasets’ descriptions are technical failures of the data logger, Internet outages, conversion work Schlemminger et al. (2022), failed radio transmissions, and problems in daylight savings time transitions Open Power System Data (2020).
@@
  • the Lower Saxony and REFIT dataset has long gaps in the measurements.
@@
3.3 Data Preprocessing
3.3.1 Handling Missing Values
3.3.2 Train-Test Split
Figure 6: Train-Test split. Time spans of each time series
3.4 Evaluated Models
A condition for the selection of TSFM was the disclosure of the training data and the provision of a pre-trained model as discussed in chapter 3.1. At the time of writing, Chronos Ansari et al. (2024), TimesFM Das et al. (2024), Sundial Liu et al. (2025), Time-MoE Shi et al. (2024), Moirai Woo et al. (2024) and LagLlama Rasul et al. (2024) fulfill both conditions. The TSFM are not be fine-tuned on the household datasets introduced for our evaluation. Only inference is used on the unseen data for achieving zero-shot predictions Xian et al. (2019). Therefore we excluded models which need finetuning e.g. the TSFM Moment Goswami et al. (2024), which forecasting head is randomly initialized. The implemention is done by following the suggested usage from the official Github repository and using the provided model weights. For LagLlama, we activate RoPe scaling as suggested in the zero-shot tutorial for input sizes longer than the context length of 32, because most of the tested input sizes are above that threshold. As the main TFS competitors for the foundation models, we used time series models with a similar Transformer-based architecture. Namely, we chose the original Transformer Encoder-Decoder implementation ("VanillaTransformer") Vaswani et al. (2017), and PatchTST Nie et al. (2023), which represents the current SOTA for household energy STLF (see Table 1). Furthermore, we chose iTransfomer Liu et al. (2024) and Temporal Fusion Transformer (TFT) Lim et al. (2021), which we identified in our literature review. The TFS Transformers are trained univariate, which can be a limiting factor for the TFS Transformers to varying degrees. However, this approach allows for the most direct comparison with the foundation models. For implementing the TFS Transformers, we used the NeuralForecast library, which is referring its model implementations to the corresponding papers.  [^1] The complete preprocessing, training and evaluation code can be found under this anonymous github repo: https://anonymous.4open.science/r/benchmarking_tsfm_household_load_forecasting-DF75/
As a baseline, we take the SeasonalAverage Hyndman and Athanasopoulos (2018) with a seasonality of 24 as we observed a high autocorrelation in the data with lags of 24. The SeasonalAverage is calculated per time series with the Pandas library with an hourly average over the input size as forecasts for the horizon. Additionally, we report as baseline reference the Naive Forecast where the last observed value is the forecast for the complete forecast horizon.
This study’s primary objective is to assess the different approaches’ general ability to adapt to new data. We vary the input size for all models to identify the influence of data context for pattern recognition. We also refrain from extensive hyperparameter tuning. We gathered over 2 million training data points, which, in combination with compute-intensive TFS Transformer training, would require a substantial amount of compute resources for hyperparameter tuning. Thus, we use the default parameters of the models as proposed in their implementation, wich are usually determined as optimal for large datasets in the corresponding papers.
A complete overview of the used models and hyperlinks to their implementation can be found in Table 3.4.
Table 5: Comparison of models, their architectures and implementations
@@ 3.4 Evaluated Models
Moirai Transformer-based Foundation Model (F) Moirai Github
Time-MoE Transformer-based Foundation Model (F) Time-MoE Github
Sundial Transformer-based Foundation Model (F) Sundial Github
@@
3.5 Training and Evaluation
We follow the approach of a time series cross-validation Bergmeir and Benítez (2012) with a calibration window. The TFS Transformers are trained on a fixed window size (365 days) of the most recent observations up to the split date. After the initial forecast, both the split date and the calibration window are moved forward by the duration of the horizon. Consequently, the training data now encompasses the most recent 365 days, starting from this updated split date. The TFS models are retrained from scratch on the updated calibration window, generating new forecasts for the next horizon period.
@@ 3.5 Training and Evaluation
Following the insights that the global is the best training approach Hertel et al. (2023), we train the TFS Transformers on all the datasets’ training splits together.
@@
For technical reasons, not all models were able to generate forecasts for every scenario. Therefore, we restricted the analysis to the subset of data for which all models provided forecasts. The resulting data reduction was only 0.2%. In total each models will be evaluated with the different setups on over 6 million forecasting points.
3.6 Challenges in Evaluating Foundation Models
While we prevent information leakage of global temporal patterns in the globally trained TFS Transformer models, there remains a special challenge in evaluating time series foundation models like Chronos Ansari et al. (2024), TimesFMDas et al. (2024) or LagLlama Rasul et al. (2024). The pre-trained models might have learned global temporal patterns (e.g., Covid-19 Prabowo et al. (2023), geopolitical crises) due to the massive amount of data used during training. Any evaluation on a dataset that is in the same date range as the training data of the foundation model could be potentially affected by such information leakage. Even cross-domain influences are possible, e.g., as weather affects household energy consumption Kang and Reiner (2022), a foundation model trained on historical weather data could theoretically lead to information leakage on a household energy dataset in the same time period.
@@ 3.6 Challenges in Evaluating Foundation Models
3.7 Metrics
The prediction results and the ground truth are normalized by subtracting the means and dividing by the standard deviation to be able to compare the time series with different demand ranges, for instance, given by the sizes of the houses and the number of occupants. The focus of the evaluation is the relative performance between models.
@@ 3.7 Metrics
$MAE_{h}=\frac{1}{h}\sum_{h=1}^{h}\left(\frac{1}{n}\sum_{i=1}^{n}|y_{h,i}-\hat{y}_{h,i}|\right)$ (1)
@@
The same logic is applied to compute the Mean Squared Error per household (MSEh). In addition, we use the adjusted p-norm error per house (APNEh), introduced by Haben et al. (2014), which is specifically designed to address the “double penalty” effect. This effect occurs when a forecast that correctly predicts the magnitude of the target (such as a peak) but is slightly displaced in time and is penalized more heavily than a constant, less informative forecast. Traditional point-wise metrics like MAE or MSE fail to account for such temporal misalignments. The adjusted p-norm error mitigates this by searching for a restricted temporal permutation of the forecast that minimizes the error according to a specified p-norm. This search is constrained by an adjustment limit window w, which defines the maximum allowable shift between the forecasted and actual time points. Following the recommendation of the authors, we use a p-norm of 4 and only adapt the adjustment window from w = 3 to w = 1 as we have hourly data instead of half-hourly data.
@@
Based on the various occurrences of zero or near-zero values, we discard other metrics, such as the mean percentage error (MAPE) and symmetric MAPE (SMAPE), as they tend to increase drastically with values near to zero Hyndman and Koehler (2006).
To ensure better comparability and to avoid imbalances between datasets, a separate ranking was established for each household across all datasets based on the (MAEh) metric. Afterwards, an average ranking was calculated across all these household rankings. This approach allows for fairer model comparisons, especially since three models could not be evaluated on the Ideal dataset and therefore had to be excluded from the analysis of that particular dataset.
4 Results
@@
Table 6: Model’s MAEh, MSEh and APNEh scores across all datasets
IDEAL Lower Saxony REFIT Southern Germany
Type Model MAEh MSEh APNEh MAEh MSEh APNEh MAEh MSEh APNEh MAEh MSEh APNEh
TSFM Chronos 0.537 1.184 2.194 0.528 1.167 2.211 0.576 1.123 2.105 0.521 0.923 1.778
Chronos-Bolt 0.513 1.010 2.073 0.490 0.931 2.019 0.528 0.880 1.932 0.478 0.731 1.616
LagLLama 0.647 1.366 2.269 0.670 1.388 2.275 0.743 1.445 2.229 0.646 1.153 1.876
Moirai 1.1 -* -* -* 1.080 3.305 4.740 1.768 >>10 >>10 1.149 4.957 6.360
Sundial -* -* -* 0.499 0.894 1.986 0.529 0.827 1.888 0.497 0.714 1.594
Time-MoE -* -* -* 0.533 0.892 1.958 0.566 0.845 1.881 0.521 0.710 1.563
TimesFM 0.514 0.998 2.063 0.491 0.929 2.026 0.530 0.885 1.939 0.478 0.722 1.627
TimesFM 2.0 0.509 1.035 2.102 0.493 0.974 2.069 0.532 0.919 1.976 0.485 0.758 1.654
TFS PatchTST 0.516 0.969 2.048 0.494 0.955 2.058 0.535 0.899 1.960 0.499 0.760 1.652
TFT 0.577 1.126 2.144 0.580 1.147 2.161 0.635 1.171 2.117 0.616 0.973 1.778
VanillaTransformer 0.549 1.075 2.124 0.556 1.098 2.141 0.603 1.113 2.097 0.573 0.896 1.744
iTransformer 0.589 1.133 2.143 0.588 1.103 2.123 0.648 1.117 2.068 0.598 0.922 1.737
Baseline Naive-Forecast 0.703 1.502 2.217 0.608 1.311 2.266 0.733 1.280 2.060 0.718 1.272 1.868
SeasonalAverage 0.609 1.208 2.128 0.572 1.102 2.071 0.577 0.982 1.965 0.539 0.854 1.696
Dataset included in TSFM pre-training data.
Table 6 shows the MAEh, MSEh and APNEh for all models across the different datasets. Comparing the different metrics, there is no single model which dominates the benchmark. But stepping back, we can see that the familiy of TSFM is outperforming the other approaches in most cases. Except for the IDEAL dataset, all best or second best results are achieved by TSFM. For the MAEh Chronos-Bolt delivers best performances across the Lower Saxony, REFIT and Southern Germany datasets, often with TimesFM following closely. The newer version TimesFM 2.0 outperformed Chronos-Bolt on the IDEAL dataset. But also the TFS Transformer PatchTST is performing good or even outperforming the other models on the IDEAL dataset in terms of MSEh. On the other datasets Time-MoE and Sundial have the best performance for the MSEh metric. The worst-performing models overall were Lag-LLama and Moirai, which could not outperform the SeasonalAverage baseline.
@@
Table 7: Rank (MAEh) and MAEh results for every input size and horizon for all datasets. Best results are in bold, second best underlined.
Type Model Input Horizon
@@
MAEh RANK MAEh RANK MAEh RANK
TSFM Chronos 24 0.55 5.29 0.55 4.82 0.57 4.63
Chronos-Bolt 24 0.53 4.13 0.54 3.94 0.56 4.10
LagLLama 24 0.60 8.59 0.78 10.86 0.82 10.77
Moirai 1.1* 24 0.75 13.50 1.98 14.00 2.56 14.00
Sundial* 24 0.53 5.17 0.56 5.23 0.56 4.88
Time-MoE* 24 0.59 9.40 0.58 7.83 0.58 7.04
TimesFM 24 0.53 3.76 0.54 3.65 0.57 5.17
TimesFM 2.0 24 0.53 4.16 0.54 3.52 0.55 3.34
TFS PatchTST 24 0.51 2.36 0.53 2.53 0.54 2.65
TFT 24 0.56 6.76 0.58 6.74 0.59 6.69
VanillaTransformer 24 0.53 3.81 0.56 5.03 0.57 5.07
iTransformer 24 0.62 9.69 0.60 8.26 0.60 7.40
Baseline Naive-Forecast 24 0.68 9.77 0.69 9.18 0.71 9.00
SeasonalAverage 24 0.64 9.73 0.67 9.70 0.71 9.62
TSFM Chronos 96 0.52 5.11 0.54 4.89 0.55 5.50
Chronos-Bolt 96 0.49 2.95 0.51 2.86 0.50 2.44
LagLLama 96 0.58 8.58 0.66 10.19 0.68 10.45
Moirai 1.1* 96 0.79 12.10 0.88 13.81 1.70 14.00
Sundial* 96 0.48 3.27 0.50 3.37 0.50 3.25
Time-MoE* 96 0.53 7.96 0.53 6.46 0.53 6.88
TimesFM 96 0.49 2.86 0.50 2.89 0.51 3.59
TimesFM 2.0 96 0.49 2.52 0.50 2.50 0.51 2.95
TFS PatchTST 96 0.50 3.88 0.51 3.77 0.52 3.73
TFT 96 0.58 8.61 0.59 8.44 0.57 7.27
VanillaTransformer 96 0.53 5.74 0.57 6.82 0.56 7.00
iTransformer 96 0.61 10.06 0.59 8.56 0.57 7.94
Baseline Naive-Forecast 96 0.68 10.19 0.69 9.80 0.71 10.05
SeasonalAverage 96 0.57 8.43 0.58 8.14 0.58 7.88
TSFM Chronos 168 0.51 4.92 0.52 5.08 0.54 5.34
Chronos-Bolt 168 0.47 2.12 0.49 2.62 0.49 2.05
LagLLama 168 0.55 7.91 0.60 9.65 0.62 9.82
Moirai 1.1* 168 1.12 11.10 1.20 11.50 0.84 13.60
Sundial* 168 0.47 3.54 0.49 3.50 0.49 3.54
Time-MoE* 168 0.51 7.12 0.51 6.13 0.52 6.38
TimesFM 168 0.48 2.63 0.49 2.85 0.50 2.89
TimesFM 2.0 168 0.47 2.22 0.49 2.18 0.49 2.60
TFS PatchTST 168 0.50 4.76 0.51 4.39 0.51 4.67
TFT 168 0.58 9.28 0.60 9.14 0.59 8.34
VanillaTransformer 168 0.53 6.73 0.57 7.41 0.57 7.17
iTransformer 168 0.61 10.47 0.57 8.16 0.58 8.29
Baseline Naive-Forecast 168 0.68 10.42 0.69 10.29 0.71 10.22
SeasonalAverage 168 0.55 7.74 0.55 7.57 0.56 7.53
IDEAL dataset excluded.
In terms of APNEh, which lies the focus on load peak prediction, Time-MoE and Sundial are slightly outperforming the other models, indicating they try more to predict load peaks, but seem to be sometimes off by a timestep. In general, the gap between TSFM and the baseline SeasonalAverage is the smallest on the APNEh metric.
A more detailed analysis how the models behave with different input sizes and horizons allows Table 7, which shows the average rank based on MAEh and the MAEh by model, input size, and horizon.
Also, in this case there is no dominant model, but a clear pattern is visible: With a short input size, the TFS Transformer PatchTST is the best model but looses its position against the TSFM when the input size increases. Especially Sundial has a slight advantage over other TSFM with input sizes longer than 24 hours. Overall, the TSFM models and also PatchTST perform better when provided with a longer input size but also the advantage of all models against the baseline decreases.
@@
Furthermore, the models’ mean error increases with longer horizons, while the performance loss remains most of the time limited. The SeasonalAverage’s performance also increases with the input size, making it a strong baseline. An exception is the VanillaTransformer which performance remains the same.
TimesFM 2.0 frequently achieves one of the lowest ranks among all models, indicating that it performs very well on many time series. However, it struggles with certain cases, which leads to a slight disadvantage in the overall MAEh.
5 Discussion
Our experiments on four datasets, namely Lower Saxony, Southern Germany, IDEAL, and REFIT, show that while the best TSFM, like TimesFM, Chronos-Bolt, Time-MoE, and Sundial outperform the best TFS Transformer, like PatchTST, not all TSFM perform equally well. This directly addresses the stated research question, “Can zero-shot TSFMs match the capabilities of state-of-the-art trained-from-scratch Transformers in forecasting household electricity load?”. Our empirical evaluation demonstrates that zero-shot TSFMs not only achieve performance on par with TFS Transformers but, in certain cases, even surpass them in the context of household electricity STLF.